If one of the root of the equation 4x2 – 15x + 4p = 0 is the square of the other then value of p is ?

Grade:12

1 Answers

Saurabh Koranglekar
5 years ago
a+a^2=15/4

solve for a
then you can solve for p
a^3 = p